In this talk I’ll discuss some of the history of Galois theory and of the problem of showing particular groups are Galois groups. A refinement of this problem has to do with prescribing how Galois groups act on various modules attached to fields. At the end of the talk I’ll describe some new work with Amy Ksir on how Galois groups act on the regular differentials of algebraic curves.